the length of a rectangle is three cm longer than twice the width and that the perimeter of the rectangle is 90 cm
Set up an equation for the perimeter involving only W, the width of the rectangle
Ive got
90=2l+2w
90=2(2w)+2(w)
90=4w+2w
90=6w
after that i dont know and im not sure if i did it right in the first place

Answer by richwmiller(17219)

What about the 3 cm?
L=3+2w
p=2w+2L
p=2w+2(3+2w)
p=2w+6+4w
p=6+6w
90=6+6w
84=6w
14=w
L=3+2w=3+28=31
p=62+28=90